# WorldAPIs: The World Is Worth How Many APIs? A Thought Experiment

### v1

@@ 4 Inducing the Action/Policy Space in the Hypothetical World
The prompts provided to LLM start with a system instruction that specifies the task requirements and the WorldObject definition. The system prompt is followed by a sequence of $k$ (we use $k=10$) in-context examples. These examples are pairs of instructions and their corresponding programs, retrieved based on the semantic similarity to a particular instruction in the given tutorial. (see Figure 5 for an abridged example).
# Use Case of squeeze
# bring the sponge to the sink and squeeze out the water
find(obj=sink_0)
while sponge_0.material_properties[’filled’] != ’empty’:
squeeze(obj=sponge_0, target=sink_0)
# Use Case of insert
# attach hose attachment to vacuum
insert(obj=hose_attachment_0, target=vacuum_0)
Listing 2: Example API use cases
@@
We perform rejection sampling on program generation. Specifically, a program that either contains syntactic errors (i.e. failed in AST parsing) or does not fully follow the instructions (i.e. does not contain comments of every numbered instruction step) will be rejected and the LLM will regenerate the program. For all the verified programs, we extract the full program and all the code snippets as API use cases and add them to the corresponding pools.
@@ 5.1 Experimental Setup
5.2 Evaluation Metrics
We evaluate the quality of the induced APIs and policies by the proposed pipeline with human evaluations and characterize the resulting action space through the length of several automatic statistics.
Human evaluation.
Since we do not have a simulator environment to realize the defined/induced APIs and unroll the induced policies, we manually evaluate the programs and APIs induced on the first 50 tutorials by each pipeline variant. We define two metrics to assess the quality of APIs and programs, which aim to approximate real execution:
Automatic evaluation.
6 Results & Analysis
@@
to mark the APIs with exact/overlapping affordance to the primitive actions in existing embodied environments (Shridhar et al., 2020; Puig et al., 2018b) and use   to mark APIs that are beyond the action space of exiting environments.
@@ 6.2 Automatic Evaluation
We illustrate in Figure 3 the expansion of the API pool while iteratively inducing on the sampled 1000 wikiHow tutorials, with filtering of infrequent APIs. At all levels of filtering, the pool of APIs demonstrates a diminishing increase in size and reaches a plateau after $\sim$ 600 steps of inductions. This indicates the induction pipeline approximates the stabilized space of frequently-evoked primitive actions to be $\sim 200-400$. Moreover, the saturation of the action space also indicates that as induction proceeds, LLMs shift from frequently inducing new APIs to largely reusing existing APIs. We provide in Figure 6 the average ratio of induced APIs at each step of the induction.
Figure 4 displays the top 50 most frequently evoked APIs in the induced action space. It can be observed that the induced APIs cover a wide variety of physical actions that are necessary for carrying out the rich space of tasks in the real world. It is worth noting that the induction is only performed on a diverse yet very small fraction ($<0.5\%$) of wikiHow tutorials. And with such limited scale yields a space of hundreds of APIs. The find suggests that the approximated $\sim 200-400$ frequent primitive actions can only serve as a proxy of the lower bound of a potentially much larger action space reflected in wikiHow.
Furthermore, we highlight with   the induced APIs whose affordances are covered/supported by primitive actions of existing embodied environments (Shridhar et al., 2020; Puig et al., 2018b). The limited coverage (9 out of 50) provides evidence of the gap between action spaces of simulations and the real world, and motivates future work on simulations with richer primitive actions.
@@ 7 Conclusion and Limitations
In this work, we attempt to approximate the size and properties of the action space of a versatile embodied agent that can carry out diverse tasks in the real world. We design a thought experiment with an induction pipeline to jointly induce primitive actions as APIs and agent policies as Pythonic programs, based on the hypothetical environments defined with wikiHow tutorials. Human and automatic evaluation verifies the usability of our induced APIs and provides an approximation of the lower bound of the realistic action space (300+). Moreover, our analysis highlights the deficiency in action diversity of existing embodied environments and motivates the development of action-rich simulations.
On the other hand, the induced action space is still noisy with a relatively high portion of redundant APIs, and it is challenging to scale up the evaluation without realizing APIs and executing policies induced by our pipeline. These limitations motivate future work in several interesting directions including reducing redundancy via self-correction (Madaan et al., 2023; Jiang et al., 2024) and evaluation in video-based embodied simulations (Du et al., 2023).
Acknowledgement
@@ A.1 Defining Objects
Listing 3: Object Definition
@@ A.2 Defining Primitive actions (APIs)
A.2.1 Base APIs
Here is the list of base APIs we defined, similar to existing simulators like (cite here)
@@ A.2.1 Base APIs
Listing 4: base APIs
A.2.2 Granularity
@@
Listing 5: Situation-based API definition
@@
Listing 6: Affordance-based API definition
@@
Listing 7: Decomposition into low-Level controls
@@
Listing 8: Decomposition at proper level of Granularity
A.2.3 State Checking and Feedback Loop
@@
Listing 9: Feedback loop

### v2

@@ 4 Inducing the Action/Policy Space in the Hypothetical World
The prompts provided to LLM start with a system instruction that specifies the task requirements and the WorldObject definition. The system prompt is followed by a sequence of $k$ (we use $k=10$) in-context examples. These examples are pairs of instructions and their corresponding programs, retrieved based on the semantic similarity to a particular instruction in the given tutorial (see Figure 5 for an abridged example).
Listing 2 Example API use cases
1# Use Case of squeeze
2# bring the sponge to the sink and squeeze out the water
3 find(obj=sink_0)
4 while sponge_0.material_properties[’filled’] != ’empty’:
5 squeeze(obj=sponge_0, target=sink_0)
6
7# Use Case of insert
8# attach hose attachment to vacuum
9insert(obj=hose_attachment_0, target=vacuum_0)
@@
We perform rejection sampling on program generation. Specifically, a program that either contains syntactic errors (i.e. failed in Abstract Syntax Tree parsing) or does not fully follow the instructions (i.e. does not contain comments of every numbered instruction step) will be rejected and the LLM will regenerate the program. For all the verified programs, we extract the full program and all the code snippets as API use cases and add them to the corresponding pools.
@@ 5.1 Experimental Setup
to mark the APIs with exact/overlapping affordance to the primitive actions in existing embodied environments (Shridhar et al. 2020; Puig et al. 2018) and use   to mark APIs that are beyond the action space of exiting environments.
5.2 Evaluation Metrics
Since our main contributions are based on the thought experiments, we focus on approximating the action space grounded in the hypothetical world (§3), instead of implementing a simulation of it. An exact execution-based evaluation is beyond the scope of our work. We thus approximate the execution-based evaluation with human evaluations and characterize the resulting action space through the lens of several automatic statistics.
Human Evaluation.
Since we do not have a simulator to execute the induced APIs and policies, we manually evaluate the quality of programs and APIs for the first 50 tutorials produced by each pipeline variant. We define two metrics to measure the quality of APIs and programs:
Automatic evaluation.
6 Results & Analysis
@@ 6.2 Automatic Evaluation
We illustrate in Figure 3 the expansion of the API pool while iteratively inducing on the sampled 1000 wikiHow tutorials, with filtering of infrequent APIs. At all levels of filtering, the pool of APIs demonstrates a diminishing increase in size and reaches a plateau after $\sim$ 600 steps of inductions. This indicates the induction pipeline approximates the stabilized space of frequently-evoked primitive actions to be $\sim 200-400$. Moreover, the saturation of the action space also indicates that as induction proceeds, LLMs shift from frequently inducing new APIs to largely reusing existing APIs. In Figure 6, we provide the average ratio of induced APIs at each step of the induction.
Figure 4 displays the top 50 most frequently evoked APIs in the induced action space. It can be observed that the induced APIs cover a wide variety of physical actions that are necessary for carrying out the rich space of tasks in the real world. It is worth noting that the induction is only performed on a diverse yet very small fraction ($<0.5\%$) of wikiHow tutorials. And with such limited scale yields a space of hundreds of APIs. The find suggests that the approximated $\sim 200-400$ frequent primitive actions can only serve as a proxy of the lower bound of a potentially much larger action space reflected in wikiHow. See Figure 7 for t-SNE-based visualization of APIs based on BERT embeddings.
Furthermore, we highlight with   the induced APIs whose affordances are covered/supported by primitive actions of existing embodied environments (Shridhar et al. 2020; Puig et al. 2018). The limited coverage (9 out of 50) provides evidence of the gap between action spaces of simulations and the real world, and motivates future work on simulations with richer primitive actions.
@@ 7 Conclusion and Limitations
Here, we attempt to approximate the size and properties of the action space of a versatile embodied agent that can carry out diverse tasks in the real world. We design a thought experiment with an induction pipeline to jointly induce primitive actions as APIs and agent policies as Pythonic programs, based on the hypothetical environments defined with wikiHow tutorials. Human and automatic evaluation verifies the usability of our induced APIs and provides an approximation of the lower bound of the realistic action space (300+). Moreover, our analysis highlights the deficiency in action diversity of existing embodied environments and motivates the development of action-rich simulations.
On the other hand, the induced action space is still noisy with a relatively high portion of redundant APIs, and it is challenging to scale up the evaluation without realizing APIs and executing policies induced by our pipeline. These limitations motivate future work in several interesting directions including reducing redundancy via self-correction (Madaan et al. 2023; Jiang et al. 2024) and evaluation in video-based embodied simulations (Du et al. 2023).
Acknowledgements
@@ A.1 Defining Objects
Listing 1: Object Definition
@@ A.2 Defining Primitive actions (APIs)
Base APIs
Here is the list of base APIs we defined, similar to existing simulators like (Puig et al. 2018; Padmakumar et al. 2021).
@@
Listing 2: base APIs
Granularity
@@
Listing 3: Situation-based API definition
@@
Listing 4: Affordance-based API definition
@@
Listing 5: Decomposition into low-Level controls
@@
Listing 6: Decomposition at proper level of Granularity
State Checking and Feedback Loop
@@
Listing 7: Feedback loop